# AlloyASG: Alloy Predicate Code Representation as a Compact Structurally Balanced Graph

### v3

@@ 6.1. Laplacian of CSBASG and Mutant as a Control Operation
At the beginning of this research, we attempted to use the Laplacian matrix to represent the matrix. However, since some self-edges exist in the scope of the Alloy ASG, and a formally defined construction of a Laplacian matrix did not include this situation, a Laplacian construction for a directed graph containing self-edges is yet to be constructed. Nonetheless, technically, the CSBASG still holds its structural balance, and a direction of inquiry of mutants as control operations in a discrete-time system still holds. There are existing works (acikmese2015spectrum, 1, 26) suggesting constructions that could apply to various kinds of graphs with self-loops, which are potentially fitting constructions to apply for the application that regards a repairing process of a pair of code segments as a control system to develop further possible methods of a modification on a software system containing multiple functional code segments to be considered.
@@ 6.2. Learning of Encoding and Alloy Code Generation
In this paper, we only attempted to ensure a zero-information-loss encoding for Alloy code segments but still relied on the map of nodes since we have a static, same-distance pre-defined encoding. Nevertheless, in theory, we can stop this reliance by ensuring no different pair of edges have the same angular signature difference.
@@
Consider a vector $\bm{\zeta}=[1\angle\theta_{1},1\angle\theta_{2},...,1\angle\theta_{n}]^{T}$ where $n$ is the number of total nodes, such that, by Lemma 1, the eigenvector of the CSBASG. We could train this vector as an embedding of the nodes in the CSBASG, giving their angular signatures. In a future scenario, such a vector could be trained with some objective function in applications such as Alloy code generation. A common approach could be an optimization problem like
$\min\sum_{i,j=1...n}\mathbb{I}((v_{i},v_{j})\in\mathcal{E})(\theta_{i}-\theta_{j})$
@@
given $\theta_{i}\neq\theta_{j}$ for any $i\neq j$. For a sub-ASG containing solely expressions or formulae with such a vector $\bm{\zeta}$, we could estimate that it could be a high probability for an incomplete ASG to have edges connecting to the most possible nodes that have a lower angular signature distance with the parent node, helping us to fill out the blanks in an incomplete code segment or generate examples within a predefined declarative environment.
6.3. Automatic Fixing by Machine Learning
Our research was motivated by a requirement to take the programming logic into the machine learning models and create a toolset that could be utilized in automatic fixing. However, due to the limitation of our dataset, few models could be fixed by a small number of mutant operations. Nonetheless, since we already have a mechanism that could output the common edges, edges to be removed, and edges to be established, we could begin with a predictive model for the probability of the existence of each edge, as suggested by some classical graph neural networks, but adapting those methods on the directed multigraph with a polynomial encoding is still posing challenges to the future research.
@@ 7. Conclusion
This paper introduced a novel graph representation of the programming language as a complex-valued multigraph with a polynomial-based adjacency matrix encoding. It evaluated its applicability on the declarative specification language Alloy. It further evaluated its compactness improvement on a dataset consisting of multiple Alloy models and the comparability of the Alloy predicates in an attempt to develop an automatic fixing framework. It points out a direction of inquiry in utilizing the representation in multiple software engineering practices utilizing Alloy and possible integration in a machine learning algorithm.

### v4

@@ 6.1. Laplacian of CSBASG and Mutant as a Control Operation
At the beginning of this research, we attempted to use a Laplacian matrix to represent the matrix. However, since some self-edges exist in the scope of the Alloy ASG, we could not as there is not currently a Laplacian construction for a directed graph containing self-edges. Nonetheless, technically, the CSBASG still holds its structural balance, and the ability to use mutants as control operations in a discrete-time system still holds. There are existing works (Acikmese, 2015; Preetha P et al., 2023) that could potentially lead to constructions that could apply to various kinds of graphs with self-loops, which are potentially fitting constructions to apply towards a repair process that uses a pair of code segments as a control system in order to explore additional methods of a modification on a software system containing multiple functional code segments under consideration.
6.2. Mitigation of Exponential Growth of Matrix Entry
Lemma 2 mentions a strict exponential bound $O(2^{pT})$to represent code segments within a repeat-free graph representation. Unfortunately, both the value of repeating nodes in the AST $T$ roughly increases linearly with the increase in code segment size, and the maximum number of nodes under a category of nodes $p$ could also end up a high constant that can vary between different code segments, all of which can prevent a universal, complete representation. Furthermore, even in our representation schema for an Alloy predicate body, as mentioned in Section 4, a rough number of $p=17$ was assigned for any technically infinitely extensible structures such as a ListExpr. While some intuitive solutions exist, like a linked or nested approach to break an infinitely extensible node into multiple nodes linked with each other, they are also costly for time and space. Therefore, the exponential growth of the adjacency matrix remains a challenge for future improvements.
6.3. Automated Repair Using Machine Learning
Our research was initially motivated by a desire to create an input format for machine learning models that would preserve the logical structure of the underlying language, as current machine learning repair techniques treat code as natural language, which limits their likelihood of generating valid Alloy formulas. Since we already have a mechanism that could output the common edges, edges to be removed, and edges to be established, we could begin with a predictive model for the probability of the existence of each edge, as suggested by some state-of-the-art graph neural networks, by adapting those methods on the directed multigraph with a polynomial encoding.
@@ 6.4. Learning of Encoding and Alloy Code Generation
In this paper, we only attempted to ensure zero-information-loss encoding for Alloy code segments, but we still relied on a map of nodes since we have a static, same-distance pre-defined encoding. Nevertheless, in theory, we can stop this reliance by ensuring no different pair of edges have the same angular signature difference.
@@
Consider a vector $\bm{\zeta}=[1\angle\theta_{1},1\angle\theta_{2},...,1\angle\theta_{n}]^{T}$ where $n$ is the number of total nodes, such that, by Lemma 1, the eigenvector of the CSBASG. We could train this vector as an embedding of the nodes in the CSBASG, giving their angular signatures. In a future scenario, such a vector could be trained with some objective function in applications such as Alloy code generation. A common approach could be an dual-annealing optimization problem like
$\min\sum_{(v_{i},v_{j},k)\in\mathcal{E}}|\theta_{i}-\theta_{j}|$
$\max\sum_{\forall k:(v_{i},v_{j},k)\notin\mathcal{E}}|\theta_{i}-\theta_{j}|$
@@
given $\theta_{i}\neq\theta_{j}$ for any $i\neq j$ and $(v_{i},v_{j},k)$ is the $k$-th edge in the multigraph form between $v_{i}$ and $v_{j}$, minimizing the signature differences between any pairs of nodes while maximizing the signature differences for the nodes that are not directly connected. For a sub-ASG containing solely expressions or formulae with such a vector $\bm{\zeta}$, we could estimate that it could be a high probability for an incomplete ASG to have edges connecting to the most possible nodes that have a lower angular signature distance with the parent node, helping us to fill out the blanks in an incomplete code segment or generate examples within a predefined declarative environment.
6.5. Possible Application on Other Languages
By construction, every programming language is built on a Context-Free Grammar, which we can utilize to create a complete construction and encoding for a CSBASG that can be used in place of an AST. However, our current completeness representation is built on the assertion that each node has limited extensibility. For real-world applications, there could be plenty of structures that either allow infinite extensions or, in a practical sense, have a value of the maximum links down from a category of node sufficiently high enough to hinder the CSBASG’s scalability severely. One way to mitigate this is to focus on building CSBASGs of small-scale, simple, and reusable code segments, such as a function body, and then use this representation to help determine its proximity to any known faulty code segments.
7. Related Work
In this section, we give an overview of the related work relevant to our planned future applications of a CSBASG for Alloy.
Automatic Repair of Alloy Models. Automatically repairing faulty Alloy models is a growing research field (Wang et al., 2018a; Gutiérrez Brida et al., 2023, 2021; Zheng et al., 2022; Cerqueira et al., 2022). ARepair, ICEBAR, and BeAFix are generating valid repair techniques that involve bounded exhaustive searches (Wang et al., 2018a; Gutiérrez Brida et al., 2023, 2021). TAR is a mutation-oriented repair technique designed specifically for Ally4Fun models (Cerqueira et al., 2022). ATR tries to find patches based on a preset number of templates (Zheng et al., 2022). While these techniques try to strike different balances in establishing a domain of patches to search through, the current state-of-the-art ATR can only correct 66% of the Alloy4Fun benchmark models. In addition to opening up the door to explore contextually aware machine learning repairs, our CSBASG could be incorporated into the backend of these existing techniques to try and improve their scalability, many of which define their search space using ASTs.
Code Generation for Alloy Models. HiGenA is a hint generator for Alloy4Fun exercises that uses historical edits to suggest a series changes a user can make to get to a correct solution (de Barros, 2023). HiGenA’s user study highlights that users only find the hints helpful when their solution is already close to the correct answer, indicating room for improvements for hint generation. Overall, while restricted to Alloy4Fun exercise, HiGenA does establish a baseline for us to compare our code generation techniques against. Similar to hint generation, there is an existing body of work that completes partial Alloy models. ASketch takes as input a partial Alloy model with holes and a test suite that outlines the expected behavior (Wang et al., 2018c). ASketch then generates substitutions for the holes (Wang et al., 2018b) and tries to find a completed model that passes all tests. However, despite several advances, ASketch still times out on the largest benchmark model (Sullivan, 2017; Jovanovic and Sullivan, 2022). While ASketch does not currently utilize ASTs, other synthesis strategies could be built utilizing our CSBASG representation and compared to ASketch.
@@ 8. Conclusion
This paper introduces a novel graph representation of a programming language as a complex-valued multigraph with a polynomial-based adjacency matrix encoding. First, we present how to create a CSBASG of the declarative specification language Alloy. Then, we evaluate the effectiveness of this representation by investigating the compactness improvements the CSBASG achieves compared to an AST for Alloy models, and the comparability CSBASG enables of Alloy predicates. In addition, we point out several future research directions that utilize this representation, including repair, code generation and improvements to the representation itself.